# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 8.1. Correlation decay
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}(|x-y|-|y-\xi_{j}|-|z-\xi_{j}|)}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz.$ (8.23)
@@
$\|\partial_{x}^{2}u_{f}\|_{L^{2}}^{2}+\|u_{f}\|_{L^{2}}^{2}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}^{2}$ (8.24)
@@
$\sup_{y\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$ $\leq C\|u_{f}\|_{H^{\frac{1}{2}+}}\leq C\|u_{f}\|_{H^{2}}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}.$ (8.25)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$ (8.26)
@@
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\frac{\lambda}{2}(|y-\xi_{j}|-|z-\xi_{j}|)}dz$
$\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz.$ (8.27)
@@
$\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz\bigg|\lesssim 1.$ (8.28)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$ (8.29)
@@
We now go back to the resolvent identity (8.17)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)-2\sum_{j=1}^{k}(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}})(x,y).$ (8.30)
If $|y-\xi_{j}|\leq\frac{1}{2}|x-y|$, then we use (8.26) to get
$|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}.$ (8.31)
@@
where
$|R_{j}(x,y)|\lesssim e^{-cd_{\varepsilon}}.$
This, along with (8.29), implies that
$|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}+e^{-c|\log(\varepsilon\log\frac{1}{\varepsilon})|}.$ (8.32)
@@
We now prove the diagonal regime (8.16). From the resolvent identity (8.30) and Lemma 8.2 (ii), we have that for any $x\in[-L_{\varepsilon},L_{\varepsilon}]$
$0\leq G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)\leq G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,x)\lesssim 1,$
uniformly in $\xi_{1},\dots,\xi_{k}$ since $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}$ and $G^{D}_{\xi_{1},\dots,\xi_{k}}$ are positive operators (and hence so are their Green functions) and $W_{j}\geq 0$. This completes the proof of Proposition 8.3.
@@
Under the separation condition $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}$, we have
@@
Recall that
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(x,y).$
Using the resolvent identity, we write
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\xi_{j}}(x,y)-G^{D}_{\xi_{j}}\Big(\sum_{\ell\neq j}W_{\ell}\Big)G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y),$ (8.33)
where
$G^{D}_{\xi_{j}}(x,y)=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(x,y).$
Again, using the single–bump resolvent identities
@@
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{j}}=G^{D}_{\xi_{j}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}.$ (8.34)
@@
$\|T_{j\ell}\|_{L^{2}\to L^{2}}\leq\|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\|W_{j}G^{D}_{\xi_{j}}W_{\ell}\|_{L^{2}\to L^{2}}\|G^{D}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}$ (8.37)
@@ 8.2. Maximum of the Gaussian process
$\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}(A_{\varepsilon}^{c})\lesssim e^{-cK^{2}|\log L_{\varepsilon}|},$ (8.38)
@@
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq Cp^{\frac{p}{2}}|x-y|^{\frac{p}{2}},$ (8.39)
uniformly in $(\xi_{1},\dots,\xi_{k})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $\overline{L}_{\varepsilon}$.
Using the fact that $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is a Gaussian measure, we have that for any finite $p\geq 2$
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq p^{\frac{p}{2}}\Big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]\Big)^{\frac{p}{2}}.$ (8.40)
@@
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]$ $=\big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(x)\big]-\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big]\big)$
$\hphantom{X}+\big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(y)v(y)\big]-\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(y)v(x)\big]\big)$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2},$ (8.41)
where
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$ $=G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}$ $=G^{D}_{\xi_{1},\dots,\xi_{k}}(y,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(y.x).$
@@
$\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)|$ $\leq c|y-z|$ (8.42)
$\sup_{y\in[-L_{\varepsilon},L_{\varepsilon}]}|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)|$ $\leq c|x-z|,$ (8.43)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and even in $L_{\varepsilon}$. By combining (8.41), (8.42) and (8.43),
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]\leq c|x-y|,$ (8.44)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]^{k}$ and $L_{\varepsilon}$. Combining (8.40) and (8.44) yields that for any finite $p\geq 2$,
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq cp^{\frac{p}{2}}|x-y|^{\frac{p}{2}},$ (8.45)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $\overline{L}_{\varepsilon}$. Therefore, we obtain the desired estimate (8.39). Once we have (8.39), the maximal behavior of the Gaussian field follows from Talagrand’s chaining argument. See [36, Proposition 5.1] or [37, Chapter 2].
@@
where $\ell$ ranges over the dyadic numbers $\ell=2^{j}K$, $j\geq 0$, so that the sum runs over $\ell=K,2K,4K,\dots$, and
$G_{\ell}=\Big\{\ell\sqrt{|\log L_{\varepsilon}|}\leq\|v\|_{L^{\infty}}<2\ell\sqrt{|\log L_{\varepsilon}|}\Big\}.$
Since $\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})$ and $\|v\|_{L^{2}}<\eta\sqrt{\log\frac{1}{\varepsilon}}$, using Hölder’s inequality and Lemma 8.5, we obtain
@@ 9. Ornstein–Uhlenbeck limit on normal space
$\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ (9.1)
@@
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$ (9.2)

### v2

@@ 8.1. Correlation decay
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$ $\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}(|x-y|-|y-\xi_{j}|-|z-\xi_{j}|)}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
(8.23) $\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz.$
@@
(8.24) $\|\partial_{x}^{2}u_{f}\|_{L^{2}}^{2}+\|u_{f}\|_{L^{2}}^{2}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}^{2}$
@@
(8.25) $\sup_{y\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$ $\leq C\|u_{f}\|_{H^{\frac{1}{2}+}}\leq C\|u_{f}\|_{H^{2}}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}.$
@@
(8.26) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$
@@
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz$ $\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\frac{\lambda}{2}(|y-\xi_{j}|-|z-\xi_{j}|)}dz$
(8.27) $\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz.$
@@
(8.28) $\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz\bigg|\lesssim 1.$
@@
(8.29) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$
@@
We now go back to the resolvent identities (8.17) and (8.18). If $|y-\xi_{j}|\leq\frac{1}{2}|x-y|$, then we use (8.26) to get
(8.30) $|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}.$
@@
where $|R_{j}(x,y)|\lesssim e^{-cd_{\varepsilon}}$. This, along with (8.29), implies that
(8.31) $|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}+e^{-c|\log(\varepsilon\log\frac{1}{\varepsilon})|}.$
@@
We now prove the diagonal regime (8.16). From the resolvent identity (8.17) and Lemma 8.2 (ii), we have that for any $x\in[-L_{\varepsilon},L_{\varepsilon}]$, $0\leq G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)\leq G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,x)\lesssim 1$, uniformly in $\xi_{1},\dots,\xi_{k}$ since $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}$ and $G^{D}_{\xi_{1},\dots,\xi_{k}}$ are positive operators and $W_{j}\geq 0$. This completes the proof of Proposition 8.3.
@@
Under $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=|\log(\varepsilon\log\frac{1}{\varepsilon})|$, we have
@@
Recall $G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$, defined in (8.1) and (8.2). Using the resolvent identity, we write
(8.32) $G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\xi_{j}}(x,y)-G^{D}_{\xi_{j}}\Big(\sum_{\ell\neq j}W_{\ell}\Big)G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y),$
where $G^{D}_{\xi_{j}}(x,y)=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(x,y)$. Again, using the single–bump resolvent identities
@@
(8.33) $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{j}}=G^{D}_{\xi_{j}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}.$
@@
(8.36) $\|T_{j\ell}\|_{L^{2}\to L^{2}}\leq\|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\|W_{j}G^{D}_{\xi_{j}}W_{\ell}\|_{L^{2}\to L^{2}}\|G^{D}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}$
@@ 8.2. Maximum of the Gaussian process
(8.37) $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}(A_{\varepsilon}^{c})\lesssim e^{-cK^{2}|\log L_{\varepsilon}|},$
@@
(8.38) $\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq Cp^{\frac{p}{2}}|x-y|^{\frac{p}{2}},$
uniformly in $(\xi_{1},\dots,\xi_{k})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $\overline{L}_{\varepsilon}$. Since $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is a Gaussian
(8.39) $\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq p^{\frac{p}{2}}\Big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]\Big)^{\frac{p}{2}}.$
@@
(8.40) $\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]=G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)+G^{D}_{\xi_{1},\dots,\xi_{k}}(y,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(y,x)$
@@
(8.41) $\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)|\leq c|y-z|$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and even in $L_{\varepsilon}$, and the same holds for the other variable. So (8.40) is bounded by c|xy|c|x-y|, thus (8.38) follows. Once we have (8.38), the maximal behavior of the Gaussian field follows from Talagrand’s chaining argument. See [36, Proposition 5.1] or [37, Chapter 2].
@@
where $\ell$ ranges over the dyadic numbers $\ell=2^{j}K$, $j\geq 0$, so that the sum runs over $\ell=K,2K,4K,\dots$, and $G_{\ell}=\big\{\ell\sqrt{|\log L_{\varepsilon}|}\leq\|v\|_{L^{\infty}}<2\ell\sqrt{|\log L_{\varepsilon}|}\big\}$. Since $\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})$ and $\|v\|_{L^{2}}<\eta\sqrt{\log\frac{1}{\varepsilon}}$, using Hölder’s inequality and Lemma 8.5, we obtain
@@ 9. Ornstein–Uhlenbeck limit on normal space
(9.1) $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$
@@
(9.2) $=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$